A current 5A passes through a straight wire in a uniform magnitude field of a flux density 2.0 T. Calculate the force per unit length exerted on the wire when it is inclined at 30° to the field.
1
Expert's answer
2020-05-15T08:41:17-0400

The Ampere's force


"F_A=BIl\sin\theta"

The force per unit length

"F_A/l=BI\sin\theta=2.0\:\rm T\times 5\: A\times \sin 30^{\circ}=5\: N."